OpenStudy (anonymous):

I need help with polynomials!!! I have watched multiple videos and read my math book, but I still don't get it. Here is the problem I need help with. y^2 represents an exponent 3x^2 + 6 -2x + 5x - 4x^2 + 9

jimthompson5910 (jim_thompson5910):

3x^2 and -4x^2 are like terms so you just add the coefficients to get 3 + (-4) = -1 this means 3x^2 + (-4x^2) = -1x^2 = -x^2

jimthompson5910 (jim_thompson5910):

you do the same for the other like terms

jimthompson5910 (jim_thompson5910):

giving out the answer is against the code of conduct and it doesn't help anyway

OpenStudy (anonymous):

you add together the like terms so: 3x^2-4x^2= -1x^2 -2x+5x=3x 6+9=15

OpenStudy (anonymous):

I don't want the answer I just need to know how to do it

OpenStudy (anonymous):

sorry jim_thompson5910 I was going to explain how I got my answer too

jimthompson5910 (jim_thompson5910):

alright, do it in the opposite order next time

OpenStudy (anonymous):

I just hit the wrong button.. sorry =(

jimthompson5910 (jim_thompson5910):

thats ok

OpenStudy (anonymous):

So would it be7x^2 + 15 +7x

jimthompson5910 (jim_thompson5910):

it would be 7x^2 IF the 4x^2 was positive but it's not

jimthompson5910 (jim_thompson5910):

same for the x terms 2x is not positive, so it's NOT 7x

OpenStudy (anonymous):

the negative sign goes with the number directly behind it so: 2x-5+7x+8 2x+7x=9x -5+8=3

jimthompson5910 (jim_thompson5910):

look back up at my first post to see how to combine the x^2 terms

OpenStudy (anonymous):

So -1^2?

jimthompson5910 (jim_thompson5910):

-1x^2 yes or just -x^2

OpenStudy (anonymous):

Ok. Thank you guys so much! I think I understand it now.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

glad to help